Merton enterprises has bonds on the market making annual payments, with 13 years to maturity, $1,000 par value, and selling for $825. At this price, the bonds yield 7 percent. What must the coupon rate be on Merton’s bonds
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, P = $825, n = 13, i = 7%

1- (1+0.07)13 825 = C* = 007 1000 T (1+0.07)13

1- 7.07)13 825 = C* -0.07 1000 (1.07)13

1- 11.07)13 825 =C* - + 414.9644 0.07

1- (1.07)13 410.0356 = C* 0.07

410.0356 = C *8.3577

C = $49.06

Coupon Rate = $49.06/$1000 = 4.91%